OpenStudy (anonymous):

-z/3 + 1/4=-3/4 what is z?? i cant figure it out

OpenStudy (anonymous):

z=3

OpenStudy (anonymous):

Isolate the variable. -z/3 + 1/4 = -3/4 -z/3 +1/4 + -1/4 = -3/4 + - 1/4 -z/3 = -1 -z/3 x 3 = -1 -z = -3 z = 3

OpenStudy (anonymous):

On the third to last step I meant to type -z/3 x 3 = -1 x 3

OpenStudy (anonymous):

thank yalll:))))) i appreciate it